# Leaf schemes and Hodge loci [^1] February 25, 2025

### v1

@@ 3.2 Hodge loci
Note that ${\sf T}:=F^{\frac{{m}}{2}}H^{m}_{\rm dR}(Y/V)$ is an algebraic bundle, however, the Hodge locus is a union of local analytic varieties $L$. Moreover, such an $L$ has a natural analytic scheme structure, that is, the natural ideal defining $L$ might not be reduced, and hence, ${\cal O}_{L}$ might have zero divisors or nilpotent elements, see Section 3.7.
@@
(Cattani-Deligne-Kaplan, [CDK95, Theorem 1.1]) The locus $L_{k}$ of Hodge cycles of self-intersection equal to $k$ is an algebraic subset of $F^{\frac{{m}}{2}}H^{m}_{\rm dR}(Y/V)$.
Let ${\mathfrak{R}}$ be a finitely generated subring of $\mathbb{C}$ such that $Y\to V$, ${\sf T}$ and $Y_{t}$ have smooth models over ${\rm Spec}({\mathfrak{R}})$ (we avoid introducing new notation for these models). For an explicit construction of ${\sf T}$ as an ${\mathfrak{R}}$-scheme see Section 3.7. Let $L$ be a local analytic Hodge locus as above. It is given by an ideal ${{\cal I}}\subset{\cal O}_{{\sf T}^{{\rm hol}}_{{\mathfrak{R}}_{\mathbb{Q}}},{t}}$. We define
(28) $\Theta_{{\sf T},L}:=\{{\sf v}\in\Theta_{\sf T}\ |\ {\sf v}({{\cal I}})\subset{{\cal I}}\},$
and we call it the ${\cal O}_{\sf T}$-module of vector fields in ${\sf T}$ tangent to $L$. One of our main motivations in the present text is the author’s not yet successful attempt to prove the following corollary of 3.1 without using it.
@@
Let $\Theta_{{\sf T},L}$ be the ${\cal O}_{{\sf T}}$-module of vector fields in ${\sf T}$ tangent to $L$. For all except a finite number of primes and all ${\sf v}\in\Theta_{{\sf T},L}$, the vector field ${\sf v}^{p}$ is also in $\Theta_{{\sf T},L}\otimes_{\mathfrak{R}}{\mathfrak{R}}_{p}$ (that is modulo $p$).
@@
Note that even though $L$ modulo $p$ might not make sense, as it is defined over complex numbers (actually ${\mathfrak{R}}_{\mathbb{Q}}$), see 3.4 and Section 3.10, the module of vector fields $\Theta_{{\sf T},L}\subset\Theta_{{\sf T}}$ is algebraic, so it makes sense to talk about modulo $p$ of this module and ${\sf v}^{p}\in\Theta_{{\sf T},L}\otimes_{\mathfrak{R}}{\mathfrak{R}}_{p}$. It says that there is a lift ${\sf w}\in\Theta_{\sf T}$ of ${\sf v}^{p}\in\Theta_{{\sf T}_{p}}$ such that it lies in $\Theta_{{\sf T},L}$. Note that a direct definition of $\Theta_{{\sf T}_{p},L}$ as in (28) does not make sense.
@@
(29) $\alpha:{{\bf T}}_{t}{\sf T}\times H^{\frac{m}{2}}(Y_{t},\Omega_{Y_{t}}^{\frac{m}{2}})\to H^{\frac{m}{2}+1}(Y_{t},\Omega_{Y_{t}}^{\frac{m}{2}-1}),\ t\in{\sf T}$
be the IVHS map (infinitesimal variation of Hodge structures in Griffiths and his coauthor’s terminology, see [CGGH83]). Here, ${{\bf T}}_{t}{\sf T}$ is the dual of the ${\mathfrak{R}}$-module $\mathfrak{m}_{{\sf T},{t}}/\mathfrak{m}_{{\sf T},{t}}^{2}$ and it is the tangent space of ${\sf T}$ at ${t}$. If ${t}$ is an ${\mathfrak{R}}$-valued point of ${\sf T}$ lying in a Hodge locus, by definition of ${\sf T}$, it comes together with a a Hodge cycle $\delta_{t}\in H^{m}(Y_{t},\mathbb{Z})$ and hence an element $\bar{t}\in H^{\frac{m}{2}}(Y_{t},\Omega_{Y_{t}}^{\frac{m}{2}})$. We denote by $L$ the Hodge locus corresponding to variations of $\delta_{t}$. The Hodge cycle $\delta_{t}$ is called general if $\alpha(\cdot,\bar{t})$ has maximal rank among all Hodge cycles. This is equivalent to say that the Hodge locus $L$ is typical in the sense of [BKU22].
@@
We have intentionally not used the language of reduction modulo a closed point of ${\rm Spec}({\mathfrak{R}})$ with residue field of characteristic $p$, in order to highlight the classical modulo $p$ manipulations. Note that since $Y_{t}$ is a projective smooth scheme over ${\rm Spec}({\mathfrak{R}})$, the cohomology groups $H^{i}(Y_{t},\Omega_{Y_{t}}^{j})$ are free ${\mathfrak{R}}$-mdoules. In 3.1 we might ask how big is $\Theta_{{\sf T},L}$. We have discussed this in Section 3.5 in the framework of leaf schemes. Since $L$ has a natural structure of an analytic scheme, our definition of ${\sf v}\in\Theta_{{\sf T},L}$ is stronger than the geometric definition: ${\sf v}({t})\in{{\bf T}}_{t}L$ for any complex point ${t}$ of ${\sf T}$. The following example might clarify the situation better. Let ${\sf T}:=\mathbb{A}^{2}_{\mathfrak{k}}$ with the coordinate system $(x,y)$ and $L$ be the subscheme of ${\sf T}$ given by $y^{2}=0$. In this case the Zariski tangent space of $L$ at each closed point is the whole tangent space of ${\sf T}$ at that point, and so, any vector field in ${\sf T}$ is tangent to $L$ in the geometric framework, but not necessarily in a scheme theoretic framework, that is, $y^{2}\mid{\sf v}(y^{2})$ is not valid in general. If ${\sf v}$ is a vector field in ${\sf T}$ and tangent to $L$ in the geometric sense then we have $\alpha({\sf v}({t}),\bar{t})=0$ for $t\in L$ but there is no reason to believe that $\alpha({\sf v}^{p}({t}),\bar{t})=0$ after taking reduction modulo $p$. This is the main reason why in 3.2 we assume that $\delta_{t}$ is a general Hodge cycle and hence $L$ is smooth. In this case both geometric and scheme theoretical definitions of tangency to $L$ are equivalent.
@@ 3.3 Leaf scheme
As in this text we care about the field or ring of definition of leaf schemes, we rewrite the definition of leaf scheme in [Mov22, Section 5.4] in the algebraic framework. It might be better for the reader to read first the definition over complex numbers before reading its algebraic/arithmetic version which involves some heavy notations due to the fact that we have to distinguish algebraic and holomorphic objects, and we have to insert the ring of definition into our notations. However, if there is no confusion, we will use simplified notations as in [Mov22, Chapter 5].
@@
Let ${\sf T}$ be an ${\mathfrak{R}}$-scheme, ${\cal F}(\Omega)$ be a foliation in ${\sf T}$ and ${t}$ be an ${\mathfrak{R}}$-valued point of ${\sf T}$. A subscheme $L$ of $({\sf T}^{{\rm hol}}_{\check{\mathfrak{R}}},{t})$ with ${\cal O}_{L}:={\cal O}_{{\sf T}^{\rm hol}_{\check{\mathfrak{R}}}}/{{\cal I}}$ is called a leaf scheme of ${\cal F}(\Omega)$ defined over $\check{\mathfrak{R}}$ if $\Omega\otimes_{{\cal O}_{{\sf T}}}{\cal O}_{{\sf T}^{{\rm hol}}_{\check{\mathfrak{R}}},{t}}$ and ${\cal O}_{{\sf T}^{{\rm hol}}_{\check{\mathfrak{R}}},{t}}\cdot d{{\cal I}}$ projected to $(\Omega^{1}_{{\sf T}}\otimes_{{\cal O}_{{\sf T}}}{\cal O}_{{\sf T}^{{\rm hol}}_{\check{\mathfrak{R}}},{t}})/{{\cal I}}\Omega^{1}_{{\sf T}}$ and regarded as ${\cal O}_{{\sf T}^{\rm hol}_{\check{\mathfrak{R}}},{t}}/{{\cal I}}$-modules are equal. In other words, $\Omega\otimes_{{\cal O}_{{\sf T}}}{\cal O}_{{\sf T}^{{\rm hol}}_{\check{\mathfrak{R}}},{t}}$ and ${\cal O}_{{\sf T}^{\rm hol}_{\check{\mathfrak{R}}},{t}}d{{\cal I}}$ are equal modulo ${{\cal I}}\Omega^{1}_{{\sf T}}$.
@@
If in the above definition ${{\cal I}}\subset{\cal O}_{{\sf T}^{\rm hol}_{\check{\mathfrak{R}}},{t}}$ (resp. ${{\cal I}}\subset{\cal O}_{{\sf T}^{\rm for}_{\check{\mathfrak{R}}},{t}}$ or ${{\cal I}}\subset{\cal O}_{{\sf T}_{\check{\mathfrak{R}}}}$) then we say that $L$ is a holomorphic leaf (resp. formal leaf or algebraic leaf) of ${\cal F}(\Omega)$ defined over $\check{\mathfrak{R}}$ and write it $L^{\rm hol}_{\check{\mathfrak{R}}}$ (resp. $L^{\rm for}_{\check{\mathfrak{R}}}$ or $L^{\rm alg}_{\check{\mathfrak{R}}}$) if it is necessary to emphasize its property of being holomorphic, formal or algebraic, and its ring of definition.
Let ${t}$ be a smooth point of ${\sf T}$, that is, there are $z_{1},z_{2},\ldots,z_{n}\in{\cal O}_{{\sf T}^{\rm hol}_{{\mathfrak{R}}_{\mathbb{Q}}},{t}}$ which generate it freely as ${\mathfrak{R}}_{\mathbb{Q}}$-algebra. We call $z=(z_{1},z_{2},\ldots,z_{n})$ a holomorphic coordinate system. By Frobenius theorem, see [Mov22, Theorem 5.8], we can choose such a coordinate system such that $\Omega\otimes_{{\cal O}_{\sf T}}{\cal O}_{{\sf T}^{\rm hol}_{{\mathfrak{R}}_{\mathbb{Q}}},{t}}$ is generated freely by $dz_{i},\ i=1,2,\ldots,k$. It turns out that the analytic scheme given by the ideal ${{\cal I}}=\langle z_{1},z_{2},\ldots,z_{k}\rangle$ is a leaf scheme of ${\cal F}(\Omega)$ and this is the classical notion of a leaf in the literature. We call it a general leaf. A general leaf is smooth (scheme theoretically) in the sense of [Mov22, Definition 5.8].
As the reader might have noticed from the definition of a general leaf, for a general definition of leaf scheme we assume that integers are invertible in the underlying ring and that is why we must use $\check{\mathfrak{R}}:={\mathfrak{R}}_{\mathbb{Q}}$.
@@
We say that a leaf scheme $L$ is algebraic if ${{\cal I}}\otimes_{{\mathfrak{R}}_{\mathbb{Q}}}\mathbb{C}$ is generated by elements in ${\cal O}_{{\sf T}_{\mathbb{C}}}$. In other words, there is a subscheme of $V$ of ${\sf T}_{\mathbb{C}}$ containing the point ${t}$ such that $L_{\mathbb{C}}$ is just $(V^{\rm hol},{t})$.
@@
We have an ideal $J\subset{\cal O}_{{\sf T}_{\mathbb{C}}}$ such that it is basically ${{\cal I}}$:
(30) ${{\cal I}}\otimes_{{\mathfrak{R}}_{\mathbb{Q}}}\mathbb{C}=J\otimes_{{\cal O}_{{\sf T}_{\mathbb{C}}}}{\cal O}_{{\sf T}_{\mathbb{C}}^{\rm hol},{t}}.$
Let $J=\langle P_{1},P_{2},\ldots,P_{s}\rangle$. In particular, we have regular algebraic functions $P_{i}$ defined in ${\sf T}_{\mathbb{C}}$ such that they define the germ of the analytic scheme $L$. After adding finitely many coefficients used in the expression of $P_{i}$’s to the quotient field ${\mathfrak{k}}$ of ${\mathfrak{R}}$, it follows that we have a finitely generated field extension ${\mathfrak{k}}\subset\tilde{\mathfrak{k}}$ such that $J$ is defined over $\tilde{\mathfrak{k}}$, and hence by abuse of notation, we write $J\subset{\cal O}_{{\sf T}_{\tilde{\mathfrak{k}}}}$. In (30) we can replace all $\mathbb{C}$’s with $\tilde{\mathfrak{k}}$ and this implies that the algebraic ideal $J$ is invariant under the Galois group ${\rm Gal}(\tilde{\mathfrak{k}}/{\mathfrak{k}})$. From this we can deduce that it is generated by elements defined over ${\mathfrak{k}}$, see [Wei62, page 19 Lemma 2] and so we can assume that there are new generators $P_{i}$’s of $J$ defined over ${\mathfrak{k}}$ and in (30) we can replace all $\mathbb{C}$’s with ${\mathfrak{k}}$. After multiplication of $P_{i}$’s by some elements of ${\mathfrak{R}}$ we get $P_{i}\in{\cal O}_{\sf T}$. The number $N\in{\mathfrak{R}}$ is the product of $N_{i}$’s attached to a set of generators $f_{i}$ of ${{\cal I}}$ such that $N_{i}f_{i}$ is in the ideal $\langle P_{1},P_{2},\ldots,P_{s}\rangle\subset{\cal O}_{{\sf T}}$. ∎
@@
Let ${\mathfrak{R}}$ be a finitely generated subring of $\mathbb{C}$ and ${\sf T}$ be an ${\mathfrak{R}}$-scheme. Let also ${\cal F}(\Omega)$ be a foliation on ${\sf T}$. If a leaf scheme is defined over ${\mathfrak{R}}$ then it is algebraic.
@@ 3.4 Leaf scheme modulo prime
(31) $\Theta_{{\sf T},L}:=\left\{{\sf v}\in\Theta_{{\sf T}}\ |\ {\sf v}({{\cal I}})\subset{{\cal I}}\right\},$
@@
Let ${\sf T}$ be an ${\mathfrak{R}}$-scheme, ${t}$ be an ${\mathfrak{R}}$-valued point of ${\sf T}$, ${\cal F}(\Omega)$ be a foliation on ${\sf T}$ and $L$ be a leaf scheme of ${\cal F}(\Omega)$ through ${t}$. If for all vector fields in ${\sf T}$ tangent to $L$, that is ${\sf v}\in\Theta_{{\sf T},L}$, and all but a finite number of primes $p$, the vector field ${\sf v}^{p}$ in ${\sf T}_{p}$ is also tangent to $L$ modulo $p$, that is, ${\sf v}^{p}\in\Theta_{{\sf T},L}\otimes_{\mathfrak{R}}{\mathfrak{R}}_{p}$, then the leaf $L$ is algebraic.
@@
(Ekedahl, Shepherd-Barron, Taylor and Luntz see [Bos01, Page 165] and [Kon]) Let ${\sf T}$ be an ${\mathfrak{R}}$-scheme and ${\cal F}$ be a non-singular foliation on ${\sf T}$. If for all vector fields ${\sf v}$ in ${\sf T}$ tangent to the leaves of ${\cal F}$ and all but a finite number of primes $p$, ${\sf v}^{p}$ is also tangent to the leaves of ${\cal F}$ then all the leaves of ${\cal F}$ are algebraic.
Note that this is a particular case of 3.2. The main evidence to this conjecture (and the next one) is due to Bost in [Bos01] in which he proves this with an extra hypothesis on the leaves (Liouville property).
@@
Let us take the leaf scheme $L$ defined over ${\mathfrak{R}}$. If ${\sf v}\in\Theta_{{\sf T},L}$ then we have ${\sf v}({{\cal I}})\subset{{\cal I}}$ and hence ${\sf v}^{p}({{\cal I}})\subset{{\cal I}}$. Since the leaf $L$ is also defined over ${\mathfrak{R}}$, it makes sense to talk about its reduction $L_{p}$ modulo $p$. In particular, a direct definition $\Theta_{{\sf T}_{p},L_{p}}$ as in (28) is possible and $\Theta_{{\sf T},L}\otimes_{\mathfrak{R}}{\mathfrak{R}}_{p}=\Theta_{{\sf T}_{p},L_{p}}$. This implies that ${\sf v}^{p}$ is tangent to the leaf $L_{p}$ in ${\sf T}_{p}$. In particular, ${\rm Sch}({\sf v}^{p}\in\Theta_{{\sf T},L}\otimes_{\mathfrak{R}}{\mathfrak{R}}_{p})\subset{\sf T}_{p}$ contains the point ${t}$. This is exactly the hypothesis of 3.2. ∎
@@
3.2 can be rewritten for foliations given by a single vector field as follows: Let ${\sf T}$ be an ${\mathfrak{R}}$-scheme, ${t}$ be an ${\mathfrak{R}}$-valued smooth point of ${\sf T}$, and ${\sf v}$ be a vector field in ${\sf T}$ with ${\sf v}({t})\not=0$. If for all but a finite number of primes $p$, ${\sf v}^{p}$ is tangent to $L$ (the collinearness scheme ${\sf v}||{\sf v}^{p}$ contains $L$), then $L$ is algebraic. In a similar way 3.3 in this case is the following: If we have ${\sf v}$ as before such that for all but a finite number of primes $p$, ${\sf v}$ is collinear with ${\sf v}^{p}$ at the point ${t}$ (the collinearness scheme ${\sf v}||{\sf v}^{p}$ contains the point ${t}$) then the solution $L$ of ${\sf v}$ through ${t}$ is algebraic. In a private communication with Daniel Litt and Joshua Lam, they explained the author that the second conjecture is wrong. Their counterexample is ${\sf v}:=zy\frac{\partial}{\partial y}+\frac{\partial}{\partial z}$ in $\mathbb{A}^{2}_{\mathbb{Z}}$. One can easily check that ${\sf v}^{n}y=P_{n}(z)y$, where $P_{n}$ can be computed recursively $P_{n+1}=P_{n}^{\prime}+zP_{n}$. For $n$ even and odd, $P_{n}$ is respectively even and odd polynomial. This implies that for $n$ odd we have $P_{n}(0)=0$ which implies that ${\sf v}^{p}(0,1)=0$ for all primes $p\not=2$. The solution through $(0,1)$ is given by $(z,e^{\frac{1}{2}z^{2}})$ which is not algebraic.  [^16] Apart from 3.2 we have also another correction to this conjecture which is the converse of 1.3.

### v2

@@ 3.2 Hodge loci
Note that ${\sf T}:=F^{\frac{{m}}{2}}H^{m}_{\rm dR}(Y/V)$ is an algebraic bundle, however, the Hodge locus is a union of local analytic varieties $L$. Moreover, such an $L$ has a natural analytic scheme structure, that is, the natural ideal defining $L$ might not be reduced, and hence, ${\mathcal{O}}_{L}$ might have zero divisors or nilpotent elements, see Section 3.7.
@@
(Cattani-Deligne-Kaplan, [cadeka, Theorem 1.1]) The locus $L_{k}$ of Hodge cycles of self-intersection equal to $k$ is an algebraic subset of $F^{\frac{{m}}{2}}H^{m}_{\rm dR}(Y/V)$.
Let $\mathfrak{R}$} be a finitely generated subring of $\mathbb{C}$ such that $Y\to V$, $\sf T$} and $Y_{t}$ have smooth models over ${\rm Spec}({\mathfrak{R}})$ (we avoid introducing new notation for these models). For an explicit construction of $\sf T$} as an $\mathfrak{R}$}-scheme see Section 3.7. Let $L$ be a local analytic Hodge locus as above. It is given by an ideal ${{\mathcal{I}}}\subset{\mathcal{O}}_{{\sf T}^{{\rm hol}}_{{\mathfrak{R}}_{\mathbb{Q}}},{t}}$. We define
(3.2) $\Theta_{{\sf T},L}:=\{{\sf v}\in\Theta_{\sf T}\ |\ {\sf v}({{\mathcal{I}}})\subset{{\mathcal{I}}}\},$
and we call it the ${\mathcal{O}}_{\sf T}$-module of vector fields in $\sf T$} tangent to $L$. One of our main motivations in the present text is the author’s not yet successful attempt to prove the following corollary of 3.1 without using it.
@@
Let $\Theta_{{\sf T},L}$ be the ${\mathcal{O}}_{{\sf T}}$-module of vector fields in $\sf T$} tangent to $L$. For all except a finite number of primes and all ${\sf v}\in\Theta_{{\sf T},L}$, the vector field ${\sf v}^{p}$ is also in $\Theta_{{\sf T},L}\otimes_{\mathfrak{R}}{\mathfrak{R}}_{p}$ (that is modulo $p$).
@@
Note that even though $L$ modulo $p$ might not make sense, as it is defined over complex numbers (actually $\mathfrak{R}$$_{\mathbb{Q}}$), see 3.4 and Section 3.10, the module of vector fields $\Theta_{{\sf T},L}\subset\Theta_{{\sf T}}$ is algebraic, so it makes sense to talk about modulo $p$ of this module and ${\sf v}^{p}\in\Theta_{{\sf T},L}\otimes_{\mathfrak{R}}{\mathfrak{R}}_{p}$. It says that there is a lift ${\sf w}\in\Theta_{\sf T}$ of ${\sf v}^{p}\in\Theta_{{\sf T}_{p}}$ such that it lies in $\Theta_{{\sf T},L}$. Note that a direct definition of $\Theta_{{\sf T}_{p},L}$ as in (3.2) does not make sense.
@@
(3.3) $\alpha:{{\bf T}}_{t}{\sf T}\times H^{\frac{m}{2}}(Y_{t},\Omega_{Y_{t}}^{\frac{m}{2}})\to H^{\frac{m}{2}+1}(Y_{t},\Omega_{Y_{t}}^{\frac{m}{2}-1}),\ t\in{\sf T}$
be the IVHS map (infinitesimal variation of Hodge structures in Griffiths and his coauthor’s terminology, see [CGGH1983]). Here, ${{\bf T}}_{t}{\sf T}$ is the dual of the $\mathfrak{R}$}-module $\mathfrak{m}_{{\sf T},{t}}/\mathfrak{m}_{{\sf T},{t}}^{2}$ and it is the tangent space of $\sf T$} at t{t}. If t{t} is an $\mathfrak{R}$}-valued point of $\sf T$} lying in a Hodge locus, by definition of $\sf T$}, it comes together with a a Hodge cycle $\delta_{t}\in H^{m}(Y_{t},\mathbb{Z})$ and hence an element $\bar{t}\in H^{\frac{m}{2}}(Y_{t},\Omega_{Y_{t}}^{\frac{m}{2}})$. We denote by $L$ the Hodge locus corresponding to variations of $\delta_{t}$. The Hodge cycle $\delta_{t}$ is called general if $\alpha(\cdot,\bar{t})$ has maximal rank among all Hodge cycles. This is equivalent to say that the Hodge locus $L$ is typical in the sense of [BKU].
@@
We have intentionally not used the language of reduction modulo a closed point of ${\rm Spec}({\mathfrak{R}})$ with residue field of characteristic $p$, in order to highlight the classical modulo $p$ manipulations. Note that since $Y_{t}$ is a projective smooth scheme over ${\rm Spec}({\mathfrak{R}})$, the cohomology groups $H^{i}(Y_{t},\Omega_{Y_{t}}^{j})$ are free $\mathfrak{R}$}-mdoules. In 3.1 we might ask how big is $\Theta_{{\sf T},L}$. We have discussed this in Section 3.5 in the framework of leaf schemes. Since $L$ has a natural structure of an analytic scheme, our definition of ${\sf v}\in\Theta_{{\sf T},L}$ is stronger than the geometric definition: ${\sf v}({t})\in{{\bf T}}_{t}L$ for any complex point t{t} of $\sf T$}. The following example might clarify the situation better. Let ${\sf T}:=\mathbb{A}^{2}_{\mathfrak{k}}$ with the coordinate system $(x,y)$ and $L$ be the subscheme of $\sf T$} given by $y^{2}=0$. In this case the Zariski tangent space of $L$ at each closed point is the whole tangent space of $\sf T$} at that point, and so, any vector field in $\sf T$} is tangent to $L$ in the geometric framework, but not necessarily in a scheme theoretic framework, that is, $y^{2}\mid{\sf v}(y^{2})$ is not valid in general. If $\sf v$} is a vector field in $\sf T$} and tangent to $L$ in the geometric sense then we have $\alpha({\sf v}({t}),\bar{t})=0$ for $t\in L$ but there is no reason to believe that $\alpha({\sf v}^{p}({t}),\bar{t})=0$ after taking reduction modulo $p$. This is the main reason why in 3.2 we assume that $\delta_{t}$ is a general Hodge cycle and hence $L$ is smooth. In this case both geometric and scheme theoretical definitions of tangency to $L$ are equivalent.
@@ 3.3 Leaf scheme
As in this text we care about the field or ring of definition of leaf schemes, we rewrite the definition of leaf scheme in [ho2020, Section 5.4] in the algebraic framework. It might be better for the reader to read first the definition over complex numbers before reading its algebraic/arithmetic version which involves some heavy notations due to the fact that we have to distinguish algebraic and holomorphic objects, and we have to insert the ring of definition into our notations. However, if there is no confusion, we will use simplified notations as in [ho2020, Chapter 5].
@@
[…]
@@
[…]
@@
We say that a leaf scheme $L$ is algebraic if $\mathcal{I}$$\otimes_{{\mathfrak{R}}_{\mathbb{Q}}}\mathbb{C}$ is generated by elements in ${\mathcal{O}}_{{\sf T}_{\mathbb{C}}}$. In other words, there is a subscheme of $V$ of $\sf T$$_{\mathbb{C}}$ containing the point t{t} such that $L_{\mathbb{C}}$ is just $(V^{\rm hol},{t})$.
@@
We have an ideal $J\subset{\mathcal{O}}_{{\sf T}_{\mathbb{C}}}$ such that it is basically $\mathcal{I}$}}:
(3.4) ${{\mathcal{I}}}\otimes_{{\mathfrak{R}}_{\mathbb{Q}}}\mathbb{C}=J\otimes_{{\mathcal{O}}_{{\sf T}_{\mathbb{C}}}}{\mathcal{O}}_{{\sf T}_{\mathbb{C}}^{\rm hol},{t}}.$
Let $J=\langle P_{1},P_{2},\ldots,P_{s}\rangle$. In particular, we have regular algebraic functions $P_{i}$ defined in $\sf T$$_{\mathbb{C}}$ such that they define the germ of the analytic scheme $L$. After adding finitely many coefficients used in the expression of $P_{i}$’s to the quotient field $\mathfrak{k}$} of $\mathfrak{R}$}, it follows that we have a finitely generated field extension ${\mathfrak{k}}\subset\tilde{\mathfrak{k}}$ such that $J$ is defined over $\tilde{\mathfrak{k}}$, and hence by abuse of notation, we write $J\subset{\mathcal{O}}_{{\sf T}_{\tilde{\mathfrak{k}}}}$. In (3.4) we can replace all $\mathbb{C}$’s with $\tilde{\mathfrak{k}}$ and this implies that the algebraic ideal $J$ is invariant under the Galois group ${\rm Gal}(\tilde{\mathfrak{k}}/{\mathfrak{k}})$. From this we can deduce that it is generated by elements defined over $\mathfrak{k}$}, see [Weil1946, page 19 Lemma 2] and so we can assume that there are new generators $P_{i}$’s of $J$ defined over $\mathfrak{k}$} and in (3.4) we can replace all $\mathbb{C}$’s with $\mathfrak{k}$}. After multiplication of $P_{i}$’s by some elements of $\mathfrak{R}$} we get $P_{i}\in{\mathcal{O}}_{\sf T}$. The number $N\in{\mathfrak{R}}$ is the product of $N_{i}$’s attached to a set of generators $f_{i}$ of $\mathcal{I}$}} such that $N_{i}f_{i}$ is in the ideal $\langle P_{1},P_{2},\ldots,P_{s}\rangle\subset{\mathcal{O}}_{{\sf T}}$. ∎
@@
Let $\mathfrak{R}$} be a finitely generated subring of $\mathbb{C}$ and $\sf T$} be an $\mathfrak{R}$}-scheme. Let also ${\mathcal{F}}(\Omega)$ be a foliation on $\sf T$}. If a leaf scheme is defined over $\mathfrak{R}$} then it is algebraic.
@@ 3.4 Leaf scheme modulo prime
(3.5) $\Theta_{{\sf T},L}:=\left\{{\sf v}\in\Theta_{{\sf T}}\ |\ {\sf v}({{\mathcal{I}}})\subset{{\mathcal{I}}}\right\},$
@@
Let $\sf T$} be an $\mathfrak{R}$}-scheme, t{t} be an $\mathfrak{R}$}-valued point of $\sf T$}, ${\mathcal{F}}(\Omega)$ be a foliation on $\sf T$} and $L$ be a leaf scheme of ${\mathcal{F}}(\Omega)$ through t{t}. If for all vector fields in $\sf T$} tangent to $L$, that is ${\sf v}\in\Theta_{{\sf T},L}$, and all but a finite number of primes $p$, the vector field ${\sf v}^{p}$ in ${\sf T}_{p}$ is also tangent to $L$ modulo $p$, that is, ${\sf v}^{p}\in\Theta_{{\sf T},L}\otimes_{\mathfrak{R}}{\mathfrak{R}}_{p}$, then the leaf $L$ is algebraic.
@@
(Ekedahl, Shepherd-Barron, Taylor and Luntz see [bost01, Page 165] and [Kontsevich2023]) Let $\sf T$} be an $\mathfrak{R}$}-scheme and $\mathcal{F}$} be a non-singular foliation on $\sf T$}. If for all vector fields $\sf v$} in $\sf T$} tangent to the leaves of $\mathcal{F}$} and all but a finite number of primes $p$, ${\sf v}^{p}$ is also tangent to the leaves of $\mathcal{F}$} then all the leaves of $\mathcal{F}$} are algebraic.
Note that this is a particular case of 3.2. The main evidence to this conjecture (and the next one) is due to Bost in [bost01] in which he proves this with an extra hypothesis on the leaves (Liouville property).
@@
Let us take the leaf scheme $L$ defined over $\mathfrak{R}$}. If ${\sf v}\in\Theta_{{\sf T},L}$ then we have ${\sf v}({{\mathcal{I}}})\subset{{\mathcal{I}}}$ and hence ${\sf v}^{p}({{\mathcal{I}}})\subset{{\mathcal{I}}}$. Since the leaf $L$ is also defined over $\mathfrak{R}$}, it makes sense to talk about its reduction $L_{p}$ modulo $p$. In particular, a direct definition $\Theta_{{\sf T}_{p},L_{p}}$ as in (3.2) is possible and $\Theta_{{\sf T},L}\otimes_{\mathfrak{R}}{\mathfrak{R}}_{p}=\Theta_{{\sf T}_{p},L_{p}}$. This implies that ${\sf v}^{p}$ is tangent to the leaf $L_{p}$ in ${\sf T}_{p}$. In particular, ${\rm Sch}({\sf v}^{p}\in\Theta_{{\sf T},L}\otimes_{\mathfrak{R}}{\mathfrak{R}}_{p})\subset{\sf T}_{p}$ contains the point t{t}. This is exactly the hypothesis of 3.2. ∎
@@
3.2 can be rewritten for foliations given by a single vector field as follows: Let $\sf T$} be an $\mathfrak{R}$}-scheme, t{t} be an $\mathfrak{R}$}-valued smooth point of $\sf T$}, and $\sf v$} be a vector field in $\sf T$} with ${\sf v}({t})\not=0$. If for all but a finite number of primes $p$, ${\sf v}^{p}$ is tangent to $L$ (the collinearness scheme ${\sf v}||{\sf v}^{p}$ contains $L$), then $L$ is algebraic. In a similar way 3.3 in this case is the following: If we have $\sf v$} as before such that for all but a finite number of primes $p$, $\sf v$} is collinear with ${\sf v}^{p}$ at the point t{t} (the collinearness scheme ${\sf v}||{\sf v}^{p}$ contains the point t{t}) then the solution $L$ of $\sf v$} through t{t} is algebraic. In a private communication with Daniel Litt and Joshua Lam, they explained the author that the second conjecture is wrong. Their counterexample is ${\sf v}:=zy\frac{\partial}{\partial y}+\frac{\partial}{\partial z}$ in $\mathbb{A}^{2}_{\mathbb{Z}}$. One can easily check that ${\sf v}^{n}y=P_{n}(z)y$, where $P_{n}$ can be computed recursively $P_{n+1}=P_{n}^{\prime}+zP_{n}$. For $n$ even and odd, $P_{n}$ is respectively even and odd polynomial. This implies that for $n$ odd we have $P_{n}(0)=0$ which implies that ${\sf v}^{p}(0,1)=0$ for all primes $p\not=2$. The solution through $(0,1)$ is given by $(z,e^{\frac{1}{2}z^{2}})$ which is not algebraic.  [^1] Apart from 3.2 we have also another correction to this conjecture which is the converse of 1.3.